# Critical and quasicritical behavior in a three-species dynamical model of semi-directed percolation

### v1

@@ Abstract
We investigate a three-species dynamical model whose dynamics naturally generate the semi-directed percolation cluster and show a non-equilibrium absorbing state phase transition from an active to inactive state. The critical threshold and exponents associated with the dynamic process are determined using Monte Carlo simulations. Critical behavior observed shows that the model belongs to the directed percolation (DP) universality class. Further, we consider the effect of spontaneous activity generation in the dynamical model. While this destroys the usual critical behaviour, we find that the dynamic susceptibility shows a maximum at a specific value of the control parameter, indicating a quasi-critical behaviour, similar to the findings in the case of DP models and DP-inspired models of neuronal activity with spontaneous activity generation. Interestingly, in the presence of spontaneous activity, we find that spatial and temporal correlations exhibit power-law decays at a value of the control parameter different from the quasi-critical threshold indicating that there are two effective thresholds in such a case, one where the response function is maximum and another where the spatial and temporal correlations show scale free behaviour.
@@ 1 Introduction
There exist many similarities and differences between the isotropic and directed models of percolation. The order parameter for both cases is the probability that a randomly chosen site belongs to/generates an infinite cluster of occupied or active sites. However, crucially, the nature of the phase transition and associated critical properties in the two models are different, and they belong to distinct universality classes. In contrast to the isotropic percolation (IP) model, directed percolation (DP) models are inherently non-equilibrium in nature due to the presence of absorbing state configurations in the dynamic process from which the system cannot escape once entered [7]. By and large, the IP and the DP universality classes form two major groups, the former characterizing many purely geometric or static processes and the latter characterizing many dynamic processes leading to absorbing states [8, 10, 11, 12].
@@
An interesting variant in this context is the semi-directed percolation (SDP) model introduced by Martin and Vannimenus [15], which occupies an intermediate position between isotropic and fully directed percolation models and is the focus of our present work. The model exhibits directionality along one of the dimensions while remaining isotropic in others. In earlier works [15, 16], the SDP clusters were generated without considering explicit dynamical processes. For instance, starting from a standard square lattice percolation configuration where sites are occupied with probability $p$, Knezevic [16] defined a semi-directed cluster as the one traced by a walk starting from an occupied site on the top row and allowing movement only in the downward and lateral directions, explicitly excluding upward motion. In Ref.[15] and [16], such SDP clusters were analysed using Renormalization Group (RG) techniques and series expansion methods, respectively, to obtain critical threshold and exponents. Despite the partial directionality, SDP model is found to belong to the same universality class as fully directed percolation.
@@
In this work, our first contribution is to cast the SDP problem in a dynamic framework. In particular, we introduce a three-species model whose dynamics naturally generate the semi-directed percolation cluster and show an absorbing state phase transition from an active to inactive state. We first verify the theoretical predictions of thresholds and exponents in [15, 16] by numerical simulations. Beyond this, the proposed dynamic framework allows us to consider the SDP problem as involving an active-absorbing phase transition happening in time and study the critical exponents associated with the dynamic process. Further, aided by the dynamical framework, we consider the effect of spontaneous activity generation in SDP. Dynamical models similar to DP with spontaneous activity generation have shown promise in modeling biological systems, such as cardiac and neuronal activity [17, 18] and in studies of forest fire dynamics [19, 20]. In such models, spontaneous activity generation corresponds to activation of neural cells or lighting of trees by an external mechanism (such as lightning in the case of a forest fire), which then can combine with the internally generated activity, aiding its spreading through the system.
@@
In the dynamical SDP model we consider, while the spontaneous activity generation destroys the usual critical behaviour, we find that the dynamic susceptibility still shows a maximum as a function of the control parameter. The behaviour is similar to that reported in the case of DP models in the presence of an external field [7, 8, 21, 22, 23, 24] and DP-like models of neuronal dynamics with spontaneous activity [25, 26, 27, 28, 29, 30]. We obtain the non-equilibrium Widom line corresponding to the quasicritical threshold for different strengths of spontaneous activity generation. Curiously, we find that spatial and temporal correlations exhibit power-law decays at a value of the control parameter different from the quasi-critical threshold. This indicates that in the presence of spontaneous activity, there are two effective thresholds, one where the response function is maximum and another where the spatial and temporal correlations show scale free behaviour.
@@ 2 Model definition
The dynamical version of the Semi-Directed Percolation (SDP) problem consists of a one-dimensional lattice in which each site can be in either of the three states denoted by 0, 11, and $2$. Here, 0 denotes the inactive and immune state, and $2$ represents the active state. State $1$ denotes a susceptible state for the spread of activity. The dynamical evolution involves the spontaneous change of states of sites and the spreading of activity through clusters of susceptible sites at each time step. These two processes are: P1) Spontaneous change of an active site to an immune one ($2\rightarrow 0$) and a change between susceptible and immune states ($1\longleftrightarrow 0$), and P2) Spreading of activity through each maximal contiguous set of susceptible (1s) sites due to contact with an active site ($1\rightarrow 2$ ). Each of the state changes in the former occurs independently and probabilistically, while the latter process occurs with probability one. Spontaneous generation of activity is considered by including the possibility of each maximal cluster of 1s turning into 2 with a probability $\epsilon$ independently in P1. Starting from an initial configuration at time $t=0$, the specific dynamical rules for the two processes are:
@@
It is clear from the rules of the dynamics that when $\epsilon=0$, the state of the system without any activity is an absorbing state. Starting from an initial configuration with the presence of 2s, activity may die out in a finite number of steps or sustain forever depending upon the value of the parameter $p$, resulting in an active-absorbing phase transition when $p$ is varied. When $\epsilon$ is non-zero, by definition, activity sustains for infinite time in an infinite lattice, thus shifting the phase transition point to the trivial value of $p=0$ (see Figure. 2 for an illustration of dynamics with non-zero $\epsilon$). However, as shown in previous studies of DP with spontaneous activity generation, we may still get behavior reminiscent of a phase transition, termed quasicritical behavior, indicated by the existence of a distinct maximum of susceptibility for non-zero $p$ [25, 26, 27, 28, 29, 30].
@@
From Figure. 1 (a) and (b), we can see that the activity spreads in space (horizontal direction in the figure) and forward in time (downward direction in the figure), but not backwards in time. The horizontal spread of activity in a cluster of 1s is limited by the inactive and immune 0 sites at the boundary of the cluster. Thus, from Figure. 1, we can infer that the spread of activity essentially traces an SDP cluster via the dynamics process defined (or the dynamical rules are defined in a way that this holds true). We can also construct such clusters in a non-dynamical way [16], where after having created a standard square lattice percolation configuration with sites being occupied with probability $p$, a semi-directed cluster is identified (see Figure. 1 (b)). The corresponding isotropic cluster is also shown in the Figure. 1 (b) to illustrate the difference between IP and SDP clusters.
With no spontaneous activity present ($\epsilon=0$), the transition from a regime where the activity persists to the one in which it ceases marks a non-equilibrium phase transition into an absorbing state. In the following, we first study the critical properties associated with the transition for the $\epsilon=0$ case by Monte-Carlo simulation techniques and thereafter, consider the scenario with spontaneous activity ($\epsilon\neq 0)$.
@@ 3.1 Threshold and critical exponents in the dynamic SDP model with $\epsilon=0$
Using the configurations starting from a single seed, an alternate way to obtain the threshold is to determine the number of active sites $N(t)$ at time $t$ as a function of $t$. Again, at the threshold, we expect a power-law behaviour of the form $N(t)\sim t^{\theta}$ with exponent $\theta$. For $p<p_{c}$, $N(t)$ decreases exponentially, and for $p>p_{c}$, $N(t)$ saturates at a constant value. In Figure. 5, we show the variation of $N(t)$ with $t$ where we verify the threshold value $p_{c}=0.6320(2)$. From the slopes of the straight lines corresponding to $p_{1}$ and $p_{2}$, we get the exponent value $\theta=0.31(2)$. The determined threshold value of the dynamical version of SDP is in good agreement with previously obtained thresholds for SDP models. In Ref. [15], a $p_{c}$ value of $0.6317(6)$ was obtained using RG techniques and in Ref. [16], a value of $0.631985(5)$ was obtained using series expansion methods.
@@ 3.2 Dynamical SDP with spontaneous activity
Spatial and temporal correlations are evaluated for nonzero $\epsilon$, and we find that the correlations exhibit power law decay at some specific value of $p$. In Figures. 12 (a) and (b), we show the equal time correlation plots $g_{\perp}(r,t)$ vs $r$ for different $p$ near the pseudocritical point for two different values of $\epsilon$. From Figure. 12 (a), we can infer that $g_{\perp}(r,t)$ nearly follows a straight line at a specific value of $p=0.611(4)$, indicating power-law decay. Note that the value of $p$ at which this happens is slightly lower than the values of $p$ at which the maximum of $\chi$ is obtained, which is $p=0.617(1)$. This difference becomes more pronounced at higher values of $\epsilon$ as seen in Figure. 12 (b) for $\epsilon=0.1$. Similar observations regarding the behaviour of the autocorrelation function $g_{\parallel}(\Delta t)$ can be made from Figure. 12 (c) and (d). In Figure. 13(a), we plot the value of $p$ at which power law decay in spatial and temporal correlations are seen along with the value at which the maximum of dynamic susceptibility is obtained for a given $\epsilon$. We can see that, for any $\epsilon$, the power law decays in spatial and temporal correlations develop at the same value of $p$. For low $\epsilon$, this $p$ value is also where the maximum of the dynamic susceptibility is obtained. However, for higher $\epsilon$, the values of $p$ at which correlations show power law nature and dynamic susceptibility shows a maximum are different as can be seen from Figure. 13(a).
Finally, the slope of the power laws in Figure. 12 gives the exponent ratios $\beta/\nu_{\perp}$ and $\beta/\nu_{\parallel}$ whose variations with $\epsilon$ are plotted in Figure. 13(a) along with that for $\epsilon=0$ for comparison. We can see that both the exponent ratios show an increasing trend with $\epsilon$ and go to roughly a constant value for larger $\epsilon$ values.

### v2

@@ Abstract
We investigate a one-dimensional three-species dynamical model whose dynamics naturally generate the semi-directed percolation cluster in time and show a non-equilibrium absorbing state phase transition from an active to inactive state. The critical threshold and exponents associated with the dynamic process are determined using Monte Carlo simulations. Critical behavior observed shows that the model belongs to the directed percolation (DP) universality class. Further, we consider the effect of spontaneous activity generation in the dynamical model. While, as expected, this destroys the usual critical behaviour, we find that the dynamic susceptibility shows a maximum at a specific value of the control parameter, indicating a quasi-critical behaviour, similar to the findings in the case of DP models and DP-inspired models of neuronal activity with spontaneous activity generation. Interestingly, in the presence of spontaneous activity, we find that spatial and temporal correlations exhibit power-law decays at a value of the control parameter different from the pseudo-threshold corresponding to the peak of the dynamic susceptibility, indicating that there are two pseudo-thresholds in such a case, one where the response function is maximum and another where the spatial and temporal correlations showscale-freee behaviour.
@@ 1 Introduction
There exist many similarities and differences between the isotropic and directed models of percolation. The order parameter for both cases is the probability that a randomly chosen site belongs to/generates an infinite cluster of occupied or active sites. However, the nature of the phase transition and associated critical properties in the two models are different. In contrast to the isotropic percolation (IP) model, directed percolation (DP) models are inherently non-equilibrium in nature due to the presence of absorbing state configurations in the dynamic process from which the system cannot escape once entered [7]. By and large, the IP and the DP universality classes form two major groups, the former characterizing many purely geometric or static processes and the latter characterizing many dynamic processes leading to absorbing states [8, 10, 11, 12].
@@
An interesting variant in this context is the semi-directed percolation (SDP) model introduced by Martin and Vannimenus [15], which occupies an intermediate position between isotropic and fully directed percolation models and is the focus of our present work. The model exhibits directionality along one of the dimensions while remaining isotropic in others. Its critical thresholds and correlation length exponent were evaluated in Ref.[15, 16] using transfer matrix and phenomenological renormalization techniques.
@@
In this work, our first contribution is to formulate the SDP problem as a one-dimensional dynamical model. In particular, we define a one-dimensional three-species model whose dynamics naturally generate the semi-directed percolation cluster in time and show an absorbing state phase transition from an active to inactive state. At first, using Monte-Carlo simulations, we determine the threshold and critical exponents, verifying numerically the earlier theoretical predictions [15, 16] and confirming that the dynamical SDP model belongs to the fully directed percolation (DP) universality class. Beyond this, the proposed dynamic framework allows us to consider the SDP problem as involving an active-absorbing phase transition happening in time and study the critical exponents associated with the dynamic process. Further, aided by the dynamical framework, we consider the effect of spontaneous activity generation in SDP. Dynamical models similar to DP with spontaneous activity generation have shown promise in modeling biological systems, such as cardiac and neuronal activity [17, 18] and in studies of forest fire dynamics [19, 20]. In such models, spontaneous activity generation corresponds to activation of neural cells or burning of trees by an external mechanism (such as lightning in the case of a forest fire), which then can combine with the internally generated activity, aiding its spreading through the system.
@@
In the dynamical SDP model we consider, while the spontaneous activity generation expectedly destroys the usual critical behaviour, we find that the dynamic susceptibility still shows a maximum as a function of the control parameter defining a pseudo-threshold [27]. Similar findings have been reported in the case of DP models in the presence of an external field [7, 8, 21, 22, 23, 24] and DP-like models of neuronal dynamics with spontaneous activity [25, 26, 27, 28, 29, 30]. We obtain the non-equilibrium Widom line corresponding to the maxima of the dynamic susceptibility for different strengths of spontaneous activity generation. Curiously, we find that spatial and temporal correlations exhibit power-law decays at a value of the control parameter different from the pseudo-threshold corresponding to the peak of the dynamic susceptibility. This indicates that in the presence of spontaneous activity, there are two pseudo-thresholds, one where the response function is maximum and another where the spatial and temporal correlations show scale free behaviour.
@@ 2 Model definition
The dynamical version of the Semi-Directed Percolation (SDP) problem consists of a one-dimensional lattice in which each site can be in either of the three states denoted by 0, 11, and $2$. Here, 0 denotes the inactive and immune state, and $2$ represents the active state. State $1$ denotes a susceptible state for the spread of activity. The dynamical evolution involves the spontaneous change of states of sites and the spreading of activity through clusters of susceptible sites at each time step. These two processes are: P1) Spontaneous change of an active site to an immune one ($2\rightarrow 0$) and a change between susceptible and immune states ($1\longleftrightarrow 0$), and P2) Spreading of activity through each maximal contiguous set of susceptible (1s) sites due to contact with an active site ($1\rightarrow 2$ ). Each of the state changes in the former occurs independently and probabilistically, while the latter process occurs with probability one. Spontaneous generation of activity is considered by including the possibility of each maximal cluster of 1s turning into 2 with a probability $\epsilon$ independently, via P1. Starting from an initial configuration at time $t=0$, the specific dynamical rules for the two processes are:
@@
$\textrm{Cluster of ~ 1s}\rightarrow 2~~~\textrm{with probability $\epsilon$ }$ (2)
@@
It is clear from the rules of the dynamics that when $\epsilon=0$, the state of the system without any activity is absorbing. Starting from an initial configuration with the presence of 2s in an infinite system, activity may die out in a finite number of time steps or sustain forever depending upon the value of the parameter $p$, resulting in an active-absorbing phase transition when $p$ is varied. Note that, in a finite system, for all $p<1$, there is always a nonzero probability that the population of 2s may eventually become extinct. When $\epsilon$ is non-zero, by definition, activity sustains for infinite time in an infinite lattice, thus shifting the phase transition point to the trivial value of $p=0$ (see Figure. 2 for an illustration of dynamics with non-zero $\epsilon$). However, as shown in previous studies of DP with spontaneous activity generation, we may still get behavior reminiscent of a phase transition, termed quasicritical behavior, indicated by the existence of a distinct maximum of susceptibility for non-zero $p$ [25, 26, 27, 28, 29, 30].
@@
From Figure. 1 (a) and (b), we can see that the activity spreads in space (horizontal direction in the figure) and forward in time (downward direction in the figure), but not backwards in time. The horizontal spread of activity in a cluster of 1s is limited by the inactive and immune 0 sites at the boundary of the cluster. Thus, from Figure. 1, we can infer that the spread of activity essentially traces an SDP cluster in time via the dynamic process defined on the one-dimensional lattice. We can also construct such clusters in a non-dynamical way, where after having created a standard square lattice percolation configuration with sites being occupied with probability $p$, a semi-directed cluster is identified (see Figure. 1 (b)). The corresponding isotropic cluster is also shown in the Figure. 1 (b) to illustrate the difference between IP and SDP clusters.
A natural question to ask is, can we define a dynamic 1D model with only two species (instead of three as done here), which will generate SDP clusters in time? It seems that this is not possible. To generate an SDP cluster in time, besides the update rules for each site with time, one must also define a cluster-spreading mechanism that determines the spatial extent of activity spreading in that time slice. This necessitates the introduction of a third species (immune) for marking the boundary of the cluster. In simple terms, to generate an SDP cluster in time, at each time step, check for any surviving active sites; if yes, then determine the extent of its spreading through the susceptible sites, the extent being constrained by the immune sites.
In our problem, With no spontaneous activity present ($\epsilon=0$), the transition from a regime where the activity persists to the one in which it ceases marks a non-equilibrium phase transition into an absorbing state. In the following, we first study the critical properties associated with the transition for the $\epsilon=0$ case by Monte-Carlo simulation techniques and thereafter, consider the scenario with spontaneous activity ($\epsilon\neq 0)$.
@@ 3.1 Threshold and critical exponents in the dynamic SDP model with $\epsilon=0$
Using the configurations starting from a single seed, an alternate way to obtain the threshold is to determine the number of active sites $N(t)$ at time $t$ as a function of $t$. Again, at the threshold, we expect a power-law behaviour of the form $N(t)\sim t^{\theta}$ with exponent $\theta$. For $p<p_{c}$, $N(t)$ decreases exponentially, and for $p>p_{c}$, $N(t)$ saturates at a constant value. In Figure. 5, we show the variation of $N(t)$ with $t$ where we verify the threshold value $p_{c}=0.6320(2)$. From the slopes of the straight lines corresponding to $p_{1}$ and $p_{2}$, we get the exponent value $\theta=0.31(2)$. The determined threshold value of the dynamical version of SDP is in good agreement with previously obtained thresholds for SDP models. In Ref. [15], a $p_{c}$ value of $0.6317(6)$ was obtained using transfer matrix and phenomenological renormalization techniques and in Ref. [16], a value of $0.631985(5)$ was obtained using an improved method.
@@ 3.2 Dynamical SDP with spontaneous activity
Spatial and temporal correlations are evaluated for nonzero $\epsilon$, and we find that the correlations exhibit power law decay at some specific value of $p$. In Figures. 12 (a) and (b), we show the equal time correlation plots $g_{\perp}(r,t)$ vs $r$ for different $p$ near the pseudo critical point for two different values of $\epsilon$. From Figure. 12 (a), we can infer that $g_{\perp}(r,t)$ nearly follows a straight line at a specific value of $p=0.611(4)$, indicating power-law decay. Note that the value of $p$ at which this happens is slightly lower than the values of $p$ at which the maximum of $\chi$ is obtained, which is $p=0.617(1)$. This difference becomes more pronounced at higher values of $\epsilon$ as seen in Figure. 12 (b) for $\epsilon=0.1$. Similar observations regarding the behaviour of the autocorrelation function $g_{\parallel}(\Delta t)$ can be made from Figure. 12 (c) and (d).
The behavior of correlations seen in Figure 12 as the control parameter $p$ is varied, is like that of correlations seen in DP and Ising-like models, i.e, a power-law with an exponential cutoff $g(r)\sim r^{-x}e^{-r/\xi}$ where $\xi$ is the correlation length, $x$ is a suitable exponent, and $r$ denotes spatial distance (or $\Delta t$ in the case of autocorrelation). Approaching the pseudo threshold from below, the correlations are well fitted by this form well, with a systematically increasing correlation length $\xi$, indicating that at the threshold, the correlations follow power-laws. Goodness-of-fit tests also prefer a power-law decay over other plausible forms like a stretched exponential at the threshold.
In Figure. 13 (a), we plot the value of $p$ at which power law decay in spatial and temporal correlations are seen along with the value at which the maximum of dynamic susceptibility is obtained for a given $\epsilon$. We can see that, for any $\epsilon$, the power law decays in spatial and temporal correlations develop at the same value of $p$. For low $\epsilon$, this $p$ value is also where the maximum of the dynamic susceptibility is obtained. However, for higher $\epsilon$, the values of $p$ at which correlations show power law nature and dynamic susceptibility shows a maximum are different as can be seen from Figure. 13 (a). This indicates that spatial and temporal correlations exhibit power-law decays at a value of the control parameter different from the pseudo-threshold corresponding to the peak of the dynamic susceptibility. Thus in the presence of spontaneous activity, there seem to be two pseudo-thresholds, one where the response function is maximum and another where the spatial and temporal correlations show scale free behaviour. We note that although power-laws are usually associated with the notion of criticality, there are instances where they can arise in non-critical situations as well [31, 32, 33, 34].
Finally, the slope of the power laws in Figure. 12 gives the exponent ratios $\beta/\nu_{\perp}$ and $\beta/\nu_{\parallel}$ whose variations with $\epsilon$ are plotted in Figure. 13 (b) along with that for $\epsilon=0$ for comparison. We can see that both the exponent ratios show an increasing trend with $\epsilon$ and go to roughly a constant value for larger $\epsilon$ values.
@@ 4 Conclusions
There are alternative ways of realizing the same dynamics; however, they also require three states. For example, in Fig. 14, we show an SDP cluster generated via a different but equivalent dynamic process to the one defined in Sec. 2.
There are similarities and differences between the model defined in the paper and the DP-inspired neuronal models [25, 26]. Both are, in fact, three-species models: The refractory period in the case of neuronal models essentially corresponds to an immune state (the site cannot be activated again during the refractory period). The difference lies in the fact that, in our problem, active sites remain active or become immune stochastically (Eq. 4). This means that the time for which activity or immunity (refractory period) sustains for a site is probabilistic. Another difference is that, in our problem, a susceptible site can directly become immune (Eq. 3), a feature usually not found in neuronal models (this is like going directly from a susceptible to a refractory state).
Figure 14: A different dynamic process on a 1D lattice which will generate SDP clusters in time. Here, a currently active site (state 2) can spread activity only if it survives to the next time step, and a site that loses activity remains immune (state 0) for exactly one time step. In a given time step, each active site attempts to spread infection to successive susceptible sites (state 1) on either side with a probability $p$, stopping at the first unsuccessful attempt.
@@
Code availability
Sample source code for our project is available on the GitHub repository.
https://github.com/jasnaCK/Semidirected-3species-model